OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 5) with a slope of 2 in standard form

OpenStudy (lgbasallote):

hint: use the formula \[\huge y - y_1 = m(x-x_1)\] where: m = slope x1 = x coordinate of the given point y1 = y coordinate of the given point does that help?